OpenStudy (anonymous):

Please help Compute Δy and dy for the given values of x and dx = Δx. (Round your answers to three decimal places.) y = 2x − x^2, x = 2, Δx = −0.6 Find: Δy and dy

OpenStudy (anonymous):

lloll

OpenStudy (anonymous):

lolololol

OpenStudy (anonymous):

trolololol

OpenStudy (anonymous):

Haha looool))) Okay, waste your time and energy guys)) I got a solution

OpenStudy (anonymous):

m'kay

OpenStudy (anonymous):

is this a differential approximation question?

OpenStudy (anonymous):

i got monster on tap, so theres a solution to the energy problem, and i have nothing but time, what a shame, im not wasting anything /:

OpenStudy (anonymous):

dominator, this person is a troll/racist.. do not help

OpenStudy (anonymous):

dominator140, ignore that guys. They hate American people. Yes, can you provide your solution?

OpenStudy (anonymous):

...no.

OpenStudy (anonymous):

Not asking you)))

OpenStudy (anonymous):

You have the solution, remember?

OpenStudy (anonymous):

I wanna compare))

OpenStudy (anonymous):

racism much lol

OpenStudy (anonymous):

but you are racist. why ask for help from people of another race?

OpenStudy (anonymous):

Another race? Not asking your help. I am Caucasian (White) and you are against White people

OpenStudy (anonymous):

No, I am white too.

OpenStudy (anonymous):

but that is irrelevant

OpenStudy (anonymous):

show me your solution. you got it right?

OpenStudy (anonymous):

y = 9547653821S13735784A77Y6375487 what'd you get?

OpenStudy (anonymous):

change in y = 4.4

OpenStudy (anonymous):

don't help it..

OpenStudy (anonymous):

y = 3r35azX^435 power of racim. Squared.

OpenStudy (anonymous):

y = 2x-x^2 dy/dx = 2-2x therefore y= 2-2(2)(-0.6) So if you actually wanted help there you go if not, Ill remember for next time I see your face.

OpenStudy (anonymous):

nazym always changes its avatar.. hiding from the truth. good job man.

OpenStudy (anonymous):

@dominator140 support racism

OpenStudy (anonymous):

dominator 140, thanks a lot) don't pay attention. I was just defending my nation from them coz they hate American people

OpenStudy (anonymous):

MathCurious, I changed it only once and that's not your business. Okay?))